OpenStudy (anonymous):

Solve using the addition and multiplication principles 3 1 2 1 _ _ _ - 4 (3x - 2) - 3 < 3

OpenStudy (anonymous):

12x - 8 -3 < 3 12x < 3 + 3 + 8 12x < 14 x < 14/12 x < 7/6

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

do u mind clarfyin

OpenStudy (anonymous):

do u follow that ok?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You need to use the distributive property and distribute the 4 to the 3x making it 12x and the 4 to the -2 making it -8

OpenStudy (anonymous):

well to be honest i dont know what you mean by the multiplication and addition principles. i know that i've got the right answer but i dont want to confuse you

OpenStudy (anonymous):

o ok thats what the instruction says... i see how u got the bottom what about the top?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

they are fractions

OpenStudy (anonymous):

yes i used the distribution principle in my first line

OpenStudy (anonymous):

3/4 1/2 and so on

OpenStudy (anonymous):

You need to get x by itself.

OpenStudy (anonymous):

Then you reduce it.

OpenStudy (anonymous):

sight up so with how jim worked it out thats final? and jim im not goin against just ask

OpenStudy (dumbcow):

write fractions more clearly...1/2 what you wrote is confusing

OpenStudy (anonymous):

It looks right. You can put this into a TI-89 or HP48/50 and have it do it for you.

OpenStudy (anonymous):

it was even more confusing but its identical to the textbook

OpenStudy (anonymous):

its wrote just as that

OpenStudy (anonymous):

i have a Ti-89

OpenStudy (anonymous):

TI-89 does all Algebra for you.

OpenStudy (anonymous):

i had to write it that way cause the parethesis

OpenStudy (anonymous):

You have to to show that you need to distribute.

OpenStudy (dumbcow):

\[\frac{3}{4}(3x-\frac{1}{2}) - \frac{2}{3} <\frac{1}{3}\]

OpenStudy (anonymous):

yes thats correct cow

OpenStudy (anonymous):

TI-89 can do that.

OpenStudy (anonymous):

OH!, then I might have been wrong.

OpenStudy (anonymous):

how? barley know how to work that calc

OpenStudy (anonymous):

I only did 4 (3x - 2) - 3 < 3

OpenStudy (anonymous):

yes i see

OpenStudy (anonymous):

Find "solve("

OpenStudy (anonymous):

Put in the problem like Cow did.

OpenStudy (anonymous):

then do ",x)

OpenStudy (anonymous):

"

OpenStudy (anonymous):

sightup i have no idea of what ur tryn to say

OpenStudy (dumbcow):

\[x <\frac{11}{18}\]

OpenStudy (anonymous):

On your TI-89...

OpenStudy (anonymous):

Another awesome utility to learn is wolframalpha.

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

cow what steps?

OpenStudy (anonymous):

ok sightup ima check it out

OpenStudy (anonymous):

Same steps, distribute, combine like terms and get x by itself. ALWAYS.

OpenStudy (anonymous):

so sightup just as u did do u think u can show work to get this sum that he got

OpenStudy (anonymous):

Let me work it out quick.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I get the same answer.

OpenStudy (anonymous):

You need to know how tho?

OpenStudy (anonymous):

i went to the sight u gave

OpenStudy (anonymous):

same answer yes but i need to see the work

OpenStudy (anonymous):

ok its like this.

OpenStudy (anonymous):

first thing like i said is distribute.

OpenStudy (anonymous):

9x/4 - 3/8 -2/3 < 1/3

OpenStudy (anonymous):

we are going to move over the - 3/8 -2/3 so it now read like so: 9x/4 < 1/3 + 3/8 + 2/3

OpenStudy (anonymous):

combine like terms.

OpenStudy (anonymous):

9x/4<11/8 is what we get

OpenStudy (anonymous):

just times it now and you get your answer.

OpenStudy (anonymous):

great thanx for showing me that!

OpenStudy (anonymous):

learn to use your TI-89 and you will never have to do algebra again.

OpenStudy (anonymous):

i need lessons um where can i get info on that?

OpenStudy (anonymous):

youtube.com... just type how to ..... ti-89

OpenStudy (anonymous):

o ok kool

OpenStudy (anonymous):

how to solve for 0 ti-89 for example

OpenStudy (anonymous):

or x

OpenStudy (anonymous):

i got to pee ttyl.

OpenStudy (anonymous):

k